In proton-beam therapy, a high-energy beam of protons is fired at a tumor. As the protons stop in the tumor, their kinetic energy breaks apart the tumor's DNA, thus killing the tumor cells. For one patient, it is desired to deposit 9.0×10−2 J of proton energy in the tumor. To create the proton beam, protons are accelerated from rest through a 1.1×104 kV potential difference.

What is the total charge of the protons that must be fired at the tumor? (Answer is NOT 5.11*1010)
Answer #1

Here ,

energy = 9 *10^-2 J

NOw, potential difference , V = 1.1 *10^7 V

let the total charge is Q

Q * V = energy

1.1 *10^7 * Q = 9 *10^-2

solving for Q

Q = 8.19 *10^-9 C

the charge of protons is 8.19 *10-9 C